# EditSplat: Multi-View Fusion and Attention-Guided Optimization for View-Consistent 3D Scene Editing with 3D Gaussian Splatting

### v1

@@ B.2 Alignment with Multi-View Information
To seamlessly incorporate multi-view information during the editing process, we extend the classifier-free guidance method. This allows the integration of multi-view fusion details into the diffusion process to facilitate multi-view consistent editing. Below is our Multi-view Fusion Guidance, an extended score estimate for multi-view aligned editing with classifier-free guidance, as specified in the main paper, Eq. 4:
@@
Guidance Interpretation. Each guidance scale (e.g., $s_{M}$, $s_{S}$, and $s_{T}$) effectively shifts the probability mass in favor of outputs that align with the corresponding conditioning. Specifically, $s_{M}$ biases the model toward multi-view fusion consistency through $p_{\theta}(h_{M}|h_{S},z_{t})$, while $s_{T}$ enhances adherence to the text prompt by aligning outputs with $p_{\theta}(h_{T}|h_{M},h_{S},z_{t})$. Lastly, $s_{S}$ leverages the implicit classifier $p_{\theta}(h_{S}|z_{t})$ to preserve the original content of the source image. This balanced guidance enables our model to achieve coherent multi-view editing results that align with structural and color details across views.
@@ C Attention-Guided Trimming (AGT) Details
Figure 10: Ablation study on pruning threshold k. The figure illustrates the training progress under different pruning thresholds k. Excessively high values of k result in the removal of too many Gaussians, hindering the editing process.
@@ C.1 Extracting Attention Map
C.2 Additional Ablation over iterations
@@
Fig. 9 highlights the contrast in edited results with and without pruning the Gaussians. Based on the instruction “Turn him into a Tolkien elf”, the attention map on the far left identifies the ear as the most prominent region, requiring significant modifications in the source scene. For optimal editing, our AGT first assigns attention weights to each source Gaussians. Then, the top 0.15% of Gaussians (1,296 in total) based on assigned weights, are pruned. The rendered image in the bottom row at iteration 0 shows that the upper part of the ear has been cleared. As a result, the ear is edited more effectively into the desired elf ear. In contrast, the top row shows that the remaining source Gaussians in the unpruned case interfere with the convergence to the elf ear.

### v2

@@ B.2 Alignment with Multi-View Information
To seamlessly incorporate multi-view information during the editing process, we extend the classifier-free guidance method. This allows the integration of multi-view fusion details into the diffusion process to facilitate multi-view consistent editing. Below is our Multi-view Fusion Guidance, an extended score estimate for multi-view aligned editing with classifier-free guidance, as specified in the main paper:
@@
Guidance Interpretation. Each guidance scale (e.g., $s_{M}$, $s_{S}$, and $s_{T}$) effectively shifts the probability mass toward outputs that align with the corresponding conditioning. For instance, $s_{M}$ biases the implicit classifier $p_{\theta}$ toward assigning higher probabilities to multi-view information, thereby ensuring consistent 2D editing across views and enhancing the quality of the resulting 3D edits (see Fig. 10). However, excessively increasing $s_{M}$ reduces the influence of the text guidance scale $s_{T}$, resulting in less pronounced editing effects. Conversely, $s_{S}$ preserves the original content from the source image, and $s_{T}$ promotes adherence to the provided text prompts. By carefully balancing these guidance scales, our model effectively achieves multi-view consistent edits that accurately reflect both structural and color details specified by textual instructions.
Figure 10: Influence of Guidance Scale $s_{M}$. Qualitative analysis of $s_{M}$ in MFG for 3D editing. Higher $s_{M}$ improves multi-view consistency, but excessively large value reduces the editing effect.
Figure 11: Influence of $w_{thres}$. Qualitative comparison of $w_{thres}$ in AGT for local editing. Lower $w_{thres}$ fails to preserve original structures, while higher values overly restrict the editing scope.
@@ C Attention-Guided Trimming (AGT) Details
Figure 12: Ablation study on pruning threshold k in AGT. The figure illustrates the training progress under different pruning thresholds k using the same source images and text prompt as in Fig. 9. Excessively high values of k result in the removal of too many Gaussians, hindering the editing process.
@@ C.1 Extracting Attention Map
C.2 Qualitative Analysis over Iterations
@@
Fig. 9 highlights the contrast in edited results with and without pruning the Gaussians. Based on the instruction “Turn him into a Tolkien elf”, the attention map on the far left identifies the ear as the most prominent region, requiring significant modifications in the source scene. For optimal editing, our AGT first assigns attention weights to each source Gaussians. Then, the top 0.15% of Gaussians (1,296 in total) based on assigned weights, are pruned. The rendered image in the bottom row at iteration 0 shows that the upper part of the ear has been cleared. As a result, the ear is edited more effectively into the desired elf ear. In contrast, the top row shows that the remaining source Gaussians in the unpruned case interfere with the convergence.
@@
C.3 Analysis of Local Editing
We provide an analysis of local editing in Attention-Guided Trimming (AGT), focusing specifically on the impact of the threshold $w_{thres}$ for semantic local editing. As illustrated in Fig. 11, varying $w_{thres}$ significantly affects the balance between preserving original scene details and effectively applying the intended semantic edits.
When the threshold $w_{thres}$ is set too low, attention-guided selection becomes excessively inclusive, causing unintended modifications beyond the target area. Consequently, critical structural elements of the original scene—such as the table in Fig. 11—fail to be properly preserved. In contrast, higher values of $w_{thres}$ significantly narrowing the scope of editing. Although this helps maintain structural integrity by protecting essential regions, it simultaneously limits the effectiveness of the semantic editing. As demonstrated in our qualitative evaluation, the edits within regions of secondary semantic importance (e.g., grass or background foliage in Fig. 11) become overly constrained, resulting in less pronounced visual changes and diminishing the overall impact of the edit.
Therefore, selecting an appropriate $w_{thres}$ is critical to achieving an optimal trade-off between preserving essential structures of the original scene and effectively implementing localized semantic editing. Our experiments reveal that intermediate values of $w_{thres}$ successfully balance these competing objectives, enabling precise and semantic localized editing.